I am studying for a test and was womdering if there was someone that could help me with this problem?
Find f^-1 (x) for f(x)= -x^3+2. Give the domain of f^-1

Answer by stanbon(57377)

Find f^-1 (x) for f(x)= -x^3+2. Give the domain of f^-1
--------------------
Interchange x and y and solve for y:
x = -y^3+2
y^3 = 2-x
y = (2-x)^(1/3)
-----------------
Domain of f^-1(x): All Real numbers.
Cheers,
Stan H.